OpenStudy (kaybabyluvzyhu16.):

WILL GET FANNED AND MEDALEDIm really confused i tried over and over please help. Which of the following would be equivalent to 9^2 ⋅ 9^6? 9 to the 8th power over 9 to the 7th power 92 ⋅ 97 9 to the 10th power over 9 to the 2nd power 90

OpenStudy (anonymousthief):

Do you remember your exponent rules? ab×ac=ab+c abac=ab−c Using these, try simplifying. So, 92×96=9?